in a certain country the heights of adult men are normally distributed with a mean of 68.3 inches and a standard deviation of 2.8 inches. The country's military requires that men have heights between 63 inches and 78 inches. determine what percentage of men are eligible for military based on height.
Answer #1

z score for 63 = (63 - 68.3)/2.8 = -1.89

z score for 78 = (78 - 68.3)/2.8 = 3.46

Hence,

Percentage of men are eligible for military based on height

= P(-1.89 < z < 3.46)

= 0.9704

= 97.04%